# Tracking the Flight: Exploring a Computational Framework for Analyzing Escape Responses in Plains Zebra (Equus quagga)

### v1

@@ Abstract
Ethological research is increasingly aided by the growing affordability and accessibility of drones, which enable the capture of high-resolution footage of animal movement at fine spatial and temporal scales. However, analyzing such footage presents the technical challenge of separating animal movement from drone motion. While non-trivial, computer vision techniques—such as image registration and Structure-from-Motion (SfM)—offer practical solutions. For conservationists, open-source tools that are user-friendly, require minimal setup, and deliver timely results are especially useful for efficient data interpretation. In this study, we evaluate three approaches: a bioimaging-based registration technique, an SfM pipeline, and a hybrid interpolation method. We apply these to a recorded escape event involving 44 plains zebras (Equus quagga), captured in a single drone video. Using the best-performing method, we extract individual trajectories and identify key behavioral patterns: increased alignment (polarization) during escape, a brief widening of spacing just before stopping, and tighter coordination near the group’s center. These insights highlight both the method’s effectiveness and its potential to scale to larger datasets, contributing to broader investigations of collective animal behavior.
@@ 1 Introduction
Animal grouping behavior has been studied since Galton’s 1871 essay on gregariousness in cattle and humans [9]. Marginal predation is widely seen as a key evolutionary force shaping social behavior and centripetal instincts [33]. In herding species such as wildebeest, group shape varies with speed: wave-like formations occur during walking, while faster movement produces columnar formations [10]. The Selfish Herd hypothesis posits that individuals move inward in a group when threatened to reduce predation risk [11, 34], a response observed across species [13, 31, 36, 8] and well-illustrated in sheepdog–sheep interactions [17].
Group living offers energetic advantages and predator detection benefits, including collective vigilance—described by the “many eyes” hypothesis—and predator confusion, which may be enhanced in striped species like zebras [34]. Plains zebra (Equus quagga), the focus of this study, form harems of 5–20 individuals that can temporarily aggregate into larger herds of 50+. These highly social animals use vocalizations, grooming, and coordinated responses to threats for cohesion and protection from predators like lions and hyenas [5, 27]. However, the specific strategies they use to coordinate escape remain largely unexamined.

### v2

@@ Abstract
Ethological research is increasingly aided by the growing affordability and accessibility of drones, which enable the capture of high-resolution footage of animal movement at fine spatial and temporal scales. However, analyzing such footage presents the technical challenge of separating animal movement from drone motion. While non-trivial, computer vision techniques, such as image registration and Structure-from-Motion (SfM), offer practical solutions. For conservationists, open-source tools that are user-friendly, require minimal setup, and deliver timely results are especially useful for efficient data interpretation. This study evaluates three approaches: a bioimaging-based registration technique, an SfM pipeline, and a hybrid interpolation method. We apply these to a recorded escape event involving 44 plains zebras (Equus quagga), captured in a single drone video. Using the best-performing method, we extract individual trajectories and identify key behavioral patterns: increased alignment (polarization) during escape, a brief widening of spacing just before stopping, and tighter coordination near the group’s center. These insights highlight the method’s effectiveness and the potential to scale to larger datasets, contributing to broader investigations of collective animal behavior.
@@ 1 Introduction
Among the evolutionary drivers of social behavior, marginal predation is widely regarded as a key force shaping both grouping tendencies and centripetal instincts—the tendency of individuals to move toward the center of a group [1]. This evolutionary pressure can be seen in the movement patterns of herding species such as wildebeest, where group shape varies with speed: wave-like formations occur during walking, while faster movement produces more linear, columnar arrangements [2]. These behavioral patterns align with the Selfish Herd hypothesis, which suggests that individuals move toward the center of a group when threatened, thereby reducing their personal risk of predation [3, 4]. This inward movement has been observed across a wide range of species [5, 6, 7, 8] and is clearly illustrated in studies of sheepdog–sheep interactions [9].
Group living provides benefits such as energetic efficiency and improved predator detection. These advantages include collective vigilance, as described by the “many eyes” hypothesis, and predator confusion, which may be especially effective in striped species like zebras [4]. Plains zebra (Equus quagga), the focus of this study, typically form harems of 5–20 individuals, which can temporarily aggregate into larger herds of over 50. These highly social animals maintain cohesion and defend against predators such as lions and hyenas through vocalizations, grooming, and coordinated responses to threats [10, 11]. However, the specific strategies zebras use to coordinate their escape behavior in response to predators remain largely unexamined.